Fill in the blanks with >,
 (–25) – (–42) _____ (– 42) – (–25)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(-25)–(-42) = -25 + 42 = 17
-42–(-25) = -42 + 25 = -17
17 > -17
Therefore (-25)–(-42) > (-42)–(-25)

Thanks